OpenStudy (stana):

Given: (2x - 3y^2)^9 Find: a.) sum of numerical coefficients b.) term that includes y^10

OpenStudy (campbell_st):

any thoughts on which you need help with..?

OpenStudy (stana):

i was wondering if there's a way to find the sum of the coefficients without having to solve them one by one

OpenStudy (adrimit):

use a binomial expansion. the only term with y^10 in it is \[(a_n)x^4*y^{10}\] \[a_n=2^{4}(-3)^{5}\frac{9!}{4!5!}\]

OpenStudy (campbell_st):

well not really, but I would just write then using combination notation and index laws maybe using summation notation \[\sum_{r=0}^{9} C_{r}^{n} (2)^{ n - r}~(-3)^r\] that would be my best guess for part A

OpenStudy (adrimit):

also, the sum of the coefficients is equal to the expanded polynomial evaluated at the point (1,1) \[p(1,1)=\left(2(1)-3(1^2)\right)^9=\left(2-3\right)^9=(-1)^9=-1\]

OpenStudy (adrimit):

another more general example is the following: find the sum of all coefficients of (x+y-2z)^3 that contain only x, z or both \[p(1,0,1)=(-1)^3=-1 \\(x+y)^3-6z(x+y)^2+12z^2(x+y)-8z^3 \\(x^3+3x^2y+3xy^2+y^3)-6z(x^2+2xy+y^2)+12z^2(x+y)-8z^3 \\eliminate\;all\;terms\;that\;contain\;y \\x^3-6zx^2+12z^2x-8z^3 \\1-6+12-8=13-14=-1\]